OpenStudy (anonymous):

Solve the IVP Problem. \[x''+4x'+4x = 1 + \delta (t-2); x(0)=x'(0)=0 \] What I have gotten the following which from the book is wrong, can someone help me? \[X(s^2+4s+4) = 1/s + e^{-2t}\]

OpenStudy (anonymous):

looks right to me

OpenStudy (anonymous):

solving with laplace atm

OpenStudy (anonymous):

Does it? The book is saying that that should turn into \[ x(t)=1/4(1-e^{-2t}-1/2te^{-2t}+u_2(t)(t-2)e^{-2(t-2)} \] Where as wolfram is giving me something different. http://www.wolframalpha.com/input/?i=inverse+laplace+transform+%281%2Fs+%2B+e%5E%28-2t%29%29%2F%28s%5E2%2B4s%2B4%29 I have not done it by hand yet though.

OpenStudy (anonymous):

actually i realized what's going on. all you did was find the laplace transform. now you need to manipulate it, by first isolating F(s) = X to \[X = \frac{ 1 + se^{-2s} }{ s(s^2 + 2)^2 }\] now find inverse transform x(t), which is the solution.

OpenStudy (anonymous):

let me know if you'd like help finding inverse transform

OpenStudy (anonymous):

Alright give me a moment and ill try it by hand.

OpenStudy (anonymous):

might actually be simpler to simplify like this: \[X = \frac{ 1 }{ s(s^2 + 2)^2 } + \frac{ e^{-2s} }{ (s^2 + 2)^2 }\] my bad

OpenStudy (anonymous):

Alright I got it and I found out what wolfram alpha was doing that make it look different. Thanks a lot!

OpenStudy (anonymous):

glad i could help :)